# Causality Is Key to Understand and Balance Multiple Goals in Trustworthy ML and Foundation Models

### v3

@@ 1 Introduction
In recent years, machine learning (ML) has made remarkable strides, driving breakthroughs in natural language processing (Achiam et al., 2023), computer vision (Brooks et al., 2024), and decision-making systems (Jia et al., 2024). These advancements have led to widespread adoption across diverse domains, including healthcare (Singhal et al., 2025), finance (Lee et al., 2024), education (Team et al., 2024), and social media (Bashiri & Kowsari, 2024), where ML models now play a crucial role in diagnostics, algorithmic trading, personalized learning, and content recommendation.
@@
Causality. One of the most influential causal frameworks is Pearl’s structural causal models (SCMs), which provide a systematic approach to reasoning about causality and integrating it into machine learning (Pearl, 2009b). This framework defines causality as the relationship between the variables where a change in one variable (the cause) directly leads to a change in another variable (the effect). It establishes a directional and often mechanistic link, distinguishing relationships arising from mere correlations.
@@
Position. Despite significant advancements in research on individual dimensions of trustworthy ML such as fairness, privacy, and explainability—there is a notable lack of efforts to integrate these dimensions into a cohesive and unified framework. Each ethical principle addresses distinct challenges, yet their interplay often involves intricate trade-offs, particularly concerning model performance metrics such as accuracy. For example, mitigating fairness-related biases may require adjustments that compromise predictive precision, while enhancing explainability can impose constraints on model complexity. We argue that systematically addressing these trade-offs is a critical step toward developing AI systems that are both ethically sound and operationally efficient. While causality has been applied to address individual challenges such as fairness or interpretability, its potential to address the intersection of these challenges has largely been overlooked (see Appendix A for a detailed review). In this paper, we argue that integrating causality into ML and foundation models offers a way to balance multiple competing objectives of trustworthy AI.
@@ 2.1 Causality for Trade-offs in Trustworthy ML
Fairness vs. Accuracy. Most of the statistical fairness literature focuses on improving fairness metrics while preserving accuracy as much as possible (Feldman et al., 2015; Wei & Niethammer, 2022; Wang et al., 2021a). However, fairness often comes at the cost of reduced accuracy, as mitigating bias may require either obscuring predictive features that also contribute to discrimination or constraining model predictions within fairness-imposed boundaries (Pinzón et al., 2022; Cooper et al., 2021; Zliobaite, 2015; Zhao & Gordon, 2022).
@@
A compelling example comes from the COMPAS dataset, where Black defendants were more likely to be classified as high-risk for recidivism. Traditional statistical debiasing approaches treat race as a direct cause of the risk score, but a causal analysis reveals that increased recidivism risk is confounded by heightened policing in predominantly Black neighborhoods. By explicitly modeling this causal structure, fairness-enhancing interventions can adjust for the effect of over-policing, ensuring that predictions reflect true recidivism risk rather than biased enforcement patterns. This results in a more accurate and fairer risk assessment (Chiappa, 2019; Zafar et al., 2017; Zhang et al., 2018).
Conflicting Notions of Fairness. Fairness in ML is often constrained by conflicting definitions and measurement approaches. Friedler et al. (2021) highlight the fundamental tension between the “what you see is what you get” and “we are all equal” worldviews—where the former accepts disparities based on observed merit, while the latter seeks to correct historical inequalities. Causal graphs can crisply formulate different notions of fairness (Nabi & Shpitser, 2018; Chen et al., 2024b), thus enabling feasible mitigation via path-specific causal effects (Avin et al., 2005).
@@ 2.2 Integrating Causality into ML
All of the above forms a causal ML cycle (Figure 1) in which ML is enhanced by causal knowledge, controlled by causal tools, and finally contributes to enriching scientific knowledge. We include a supplementary introduction to causality and causal ML in Appendices C and D.
@@ 3 Causality for Trustworthy Foundation Models
Foundation models, including state-of-the-art multimodal systems like Large Language Models (LLMs) and vision-language models, have demonstrated exceptional capabilities across diverse tasks (Achiam et al., 2023; Team et al., 2023; Radford et al., 2023; Brooks et al., 2024). However, their reliability remains a concern due to issues like spurious correlations, hallucinations, and unequal representation. Trade-offs and causality in trustworthy foundation models remain underexplored as compared to traditional ML. In this section, we explore the potential for causality to improve fairness, explainability, privacy, and robustness in foundation models following slightly different taxonomy than in the previous section due to their unique challenges.
@@ 3.1 Dimensions of Trustworthy Foundation Models
Fairness vs. Accuracy. Causal frameworks have become integral to fairness interventions in LLMs by identifying and mitigating pathways that lead to unfair predictions (Madhavan et al., 2023a; Cotta & Maddison, 2024). Counterfactual fairness ensures that sensitive attributes (e.g., gender, race) do not causally influence outcomes. For example, in job recommendation systems, counterfactual fairness guarantees identical recommendations for equally qualified candidates regardless of their gender (Madhavan et al., 2023a). Methods like causal disentanglement isolate sensitive features from output-relevant causal factors, ensuring that spurious correlations, such as gender biases in job roles, do not propagate through the model (Zhou et al., 2023a; Chen et al., 2024a). SCMs further enable fairness-aware fine-tuning by disentangling causal effects. However, striving for diversity has been shown to introduce non-factual output in text-to-image models. In early 2024, Google’s AI tool, Gemini, faced criticism for generating historically inaccurate images, such as depicting America’s Founding Fathers as Black individuals and Nazis as racially diverse (Vincent, 2024). Here, causality could help distinguish historically impossible scenarios from desirable diversity, ensuring both fairness and factual integrity in AI-generated content. Mode collapse is another foundation model-specific fairness issue where models generate overly generic outputs, reducing diversity and disproportionately omitting minority group representations. Causal modeling can potentially help preserve minority information by explicitly capturing causal relationships, preventing spurious correlations from erasing underrepresented patterns.
Robustness vs. Accuracy. Causal frameworks address robustness by training models to rely on invariant causal relationships while penalizing reliance on dataset-specific spurious features (Wu et al., 2024). For instance, instead of associating ”doctor” with ”male,” causal invariance enforces reliance on task-relevant features like medical terminology (Zhou et al., 2023a). Causal regularization further discourages attention to non-causal patterns during inference achieving better accuracy and robustness.
@@ 3.2 Integrating Causality in Foundation Models
Pre-training: Causal Representation Learning. By disentangling causal factors from non-causal ones, models can learn representations that separate meaningful causal features from irrelevant associations. Techniques such as causal embedding methods (Rajendran et al., 2024; Jiang et al., 2024), which can use training data annotated with causal labels, can guide models in identifying and prioritizing true causal relationships. This has been shown to reduce reliance on spurious correlations, such as gender-biased occupational associations (Zhou et al., 2023b).
@@
Post-training: Fine-tuning. Fine-tuning on datasets specifically designed to highlight causal reasoning (e.g., datasets emphasizing cause-effect linguistic patterns) ensures that models learn causal-invariant patterns. Further, counterfactual data samples can also improve the fine-tuning. Synthetic counterfactual examples improve the model’s robustness to spurious correlations, similar to pre-training, but with better sample size efficiency. Frameworks like DISCO (Chen et al., 2022) generate diverse counterfactuals during fine-tuning to enhance OOD generalization for downstream tasks. Causally Fair Language Models (CFL) (Madhavan et al., 2023b) use SCM-based regularization to detoxify outputs or enforce demographically neutral generation during post-training.  Wang & Culotta (2020) use causal reasoning to separate genuine from spurious correlations by computing controlled direct effects, ensuring robust performance.
@@ Acknowledgements
Appendix A Related work
The literature on trade-offs in ethical AI emphasizes the inherent tensions and competing objectives involved in designing and deploying AI systems that align with ethical principles (Mbiazi et al., 2023). The works of Sanderson et al. (2024); Whittlestone et al. (2019); Katirai & Nagato (2024) explore frameworks for balancing fairness, accuracy, and other conflicting priorities. Kemmerzell & Schreiner (2024) explore trade-offs between robustness, accuracy, fairness, and privacy, and suggest data augmentation techniques that minimize the trade-offs.
The surveys on the use of causality for trustworthy AI (Ganguly et al., 2023; Rawal et al., 2024; Liu et al., 2023) provide a comprehensive overview of the existing use cases. However, they do not explicitly discuss the need for causality and do not focus on the role of causality in alleviating tensions in trustworthy AI. Notably, only Ganguly et al. (2023) overview the use of causality in privacy and exclusively focuses on the adversarial robustness through generalization. Discussions on causality for ethical AI focusing on challenges and applications are provided in (Rahmattalabi & Xiang, 2022; Vallverdú, 2024).
Several works discuss the benefits of use of causality for one or two of the aspects of trustworthy AI. Discussion on the need for causality for fairness can be found in the work of (Binkytė et al., 2022; Plecko & Bareinboim, 2024; Makhlouf et al., 2024). The study by (Wang et al., 2021b; Ehyaei et al., 2023) explored causality to enhance fairness and robustness.
Appendix B Causality. Frameworks and Definitions
The field of statistical causality encompasses a diverse range of theories and approaches that often complement or compete with each other, rather than forming a unified framework. Researchers have likened the current state of statistical causality to ”probability theory before Kolmogorov” (Dawid, 2015). In practice, the application of statistical causality typically involves combining tools and methods from multiple frameworks. This section provides an overview of the existing landscape, highlighting key theories and definitions. Most approaches conceptualize causation either as a relationship revealed through linear regression, grounded in the notion of real or hypothetical interventions, or requiring a mechanistic understanding of the underlying processes (Berzuini et al., 2012). In this work, we primarily rely on the structural probabilistic models framework (Pearl, 2009a) and the potential outcomes framework (Rubin, 2005). Below, we provide an overview of these frameworks and briefly touch on other approaches to causality. For technical definitions of relevant causal concepts, refer to the Technical Preliminaries C.
B.1 Potential Outcome Framework
The potential outcomes framework is one of the earliest formal theories of causal inference (Sjölander, 2012). It defines causal effects as the difference in potential outcomes under different levels of exposure or treatment (Rubin, 2005). This framework uses the language of potential outcomes to express causal effects in terms of joint distributions of potential outcomes represented as random variables. Causal assumptions in this framework are encoded as constraints on these distributions (Shpitser, 2012).
Potential outcomes can be categorized as factual (representing what actually occurred) or counterfactual (representing what would have occurred under different conditions). For example, if an individual took a medication and recovered, the factual outcome is ”recovery,” while the counterfactual outcome represents what would have happened if the medication had not been taken. Since counterfactual outcomes are inherently unobservable for an individual, estimating subject-specific causal effects is often impractical (Sjölander, 2012).
At the population level, however, counterfactual outcomes and causal effects can be estimated. Population-level causal effects contrast outcomes when everyone receives a treatment versus when no one does. Although only factual outcomes are observed, randomization allows for causal effect estimation under the Stable Unit Treatment Value Assumption (SUTVA) (Sjölander, 2012). Randomization ensures that potential outcomes are statistically independent of exposure, enabling identification of causal effects (Rubin, 2005). These principles are formally established in the literature (Rubin, 2005; Sjölander, 2012).
While the potential outcomes framework is widely used, it has limitations. Pearl has critiqued the framework for not providing systematic guidelines on which covariates to include for adjustment (Pearl, 1988). He warns that including all available covariates may inadvertently increase bias, highlighting the need for caution when selecting adjustment variables.
B.2 Non-Parametric Structural Models (NPSEM)
The framework proposed by Pearl (Pearl, 2009a) is often celebrated for its coherence and robust formal foundations (Dawid, 2010). Pearl integrates principles from agency causality (focused on interventions), probabilistic graphical models (Dawid, 2010), and counterfactual reasoning (Sjölander, 2012). His approach balances the probabilistic view of causality from Bayesian models and the deterministic view from structural equation models (SEMs) common in econometrics and social sciences (Pearl, 2009a).
The NPSEM framework represents causal relationships using directed acyclic graphs (DAGs). A DAG $\mathcal{G}=(\mathbf{V},\mathcal{E})$ consists of a set of variables $\mathbf{V}$ and directed edges $\mathcal{E}$ that encode causal dependencies. The structure ensures no cycles are formed. DAGs connect causal structure with joint probability distributions via the Markov condition, which states that each variable is conditionally independent of its non-descendants given its parents.
DAGs not only capture conditional independence but also distinguish causal from non-causal data-generating processes. If a variable $Y$ has an incoming edge from $X$, $X$ is a direct cause of $Y$. Indirect causation is mediated by intermediate variables; for instance, if $X$ influences $Y$ via $Z$, then $Z$ is a mediator. The NPSEM framework also provides criteria for determining the identifiability of causal quantities from observational data (Pearl, 2009a), making it a powerful tool for causal inference (Shpitser, 2012).
B.3 Alternative Approaches
The sufficient cause framework views causation as a set of sufficient conditions leading to an event (VanderWeele, 2012; Mackie, 1965). Unlike the potential outcomes approach, which emphasizes causes, this framework focuses on effects (VanderWeele, 2012). Pearl extends this by proposing probabilistic notions of necessity and sufficiency (Pearl, 2009a).
The decision-theoretic approach incorporates stochastic counterfactuals to facilitate inference transportability between observational and experimental settings (Berzuini et al., 2012). This approach relaxes strong assumptions often required by potential outcomes (Dawid, 2015).
Finally, structural equation models (SEMs), rooted in deterministic relationships expressed through structural linear equations, remain widely used but are limited by their parametric assumptions and inability to model complex, non-linear causal relationships (Wright, 1921).
Appendix C Causality: Technical Preliminaries
C.1 Causal Structures
Variables are represented by capital letters (e.g., $X$, $Y$), while specific values of variables are indicated using lowercase letters (e.g., $A=a$, $W=w$). Sets of variables and their values are denoted by bold capital letters (e.g., $\mathbf{V}$) and bold lowercase letters (e.g., $\mathbf{v}$), respectively.
A causal graph, denoted as $\mathcal{G}=(\mathbf{V},\mathcal{E})$, is a Directed Acyclic Graph (DAG) consisting of a set of variables or nodes $\mathbf{V}$ and edges $\mathcal{E}$. Each edge $X\to Y$ signifies a causal relationship, meaning changes in $X$ directly influence $Y$. Importantly, altering $X$ impacts $Y$, but modifying $Y$ does not affect $X$.
Causal graphs include three foundational structures: mediators, confounders, and colliders (Pearl, 2009b), as illustrated in Figure 3:
  • Mediator: A variable $W$ (Figure 3(a)) mediates the effect of $X$ on $Y$. For instance, $X\to W\to Y$ shows $X$’s influence on $Y$ through $W$. Mediators are also called chain structures.
  • Confounder: A variable $C$ (Figure 3(b)) is a common cause of $X$ and $Y$, resulting in a non-causal correlation between them. While $X$ and $Y$ are correlated in this structure, $X$ does not directly cause $Y$.
  • Collider: A variable $Z$ (Figure 3(c)) is influenced by $X$ and $Y$. Unlike the other structures, $X$ and $Y$ are uncorrelated unless conditioned on $Z$. Colliders are also known as v-structures.
XYW
(a) Mediator
XYC
(b) Confounder
XYZ
(c) Collider
Figure 3: Basic structures of causal graphs.
Mediation Analysis
Causal relationships often involve multiple pathways, requiring mediation analysis to distinguish between them. For example, the causal effect between $X$ and $Y$ can be decomposed into:
  • Direct effect: The path $X\to Y$.
  • Indirect effects: Paths such as $X\to R\to Y$ and $X\to E\to Y$.
  • Path-specific effects: Effects through a specific path, such as $X\to E\to Y$.
This decomposition is critical for fairness. A direct effect of $X$ on $Y$ is typically considered unfair when $X$ is a sensitive attribute (e.g., gender or race). In contrast, indirect effects may be fair or unfair, depending on the mediator. For example:
  • An indirect effect through a discriminatory variable ($R$) is unfair.
  • An indirect effect through an acceptable explanatory variable ($E$) is considered fair.
A variable is deemed a proxy (e.g., $R$) if it serves as a substitute for $X$ and leads to the same discriminatory outcome. Determining whether a variable is a proxy or an acceptable mediator often requires domain expertise.
C.2 Causal Fairness Notions
Causal fairness aims to ensure that sensitive attributes, such as race or gender, do not unfairly influence outcomes. Below, we describe key causal fairness notions and their formal definitions.
C.2.1 Total Effect (TE)
Total Effect (TE) (Pearl, 2009b) is a causal fairness notion that quantifies the overall effect of a sensitive attribute $X$ on an outcome $Y$. Formally, TE is defined as:
$TE_{x_{1},x_{0}}(y)=\mathbb{P}(Y=y\mid do(X=x_{1}))-\mathbb{P}(Y=y\mid do(X=x_{0})),$ (1)
where $do(X=x)$ denotes an intervention that sets $X$ to $x$. TE measures the causal impact of changing $X$ from $x_{0}$ to $x_{1}$ on $Y$ across all causal paths connecting $X$ to $Y$.
C.2.2 Mediation Analysis: NDE, NIE, and PSE
Mediation analysis decomposes the causal effect of $X$ on $Y$ into direct and indirect effects. This is essential for identifying the pathways through which $X$ influences $Y$.
Natural Direct Effect (NDE) (Pearl, 2001): The NDE quantifies the direct effect of $X$ on $Y$, bypassing any mediators. For a binary variable $X$ with values $x_{0}$ and $x_{1}$, the NDE is:
$NDE_{x_{1},x_{0}}(y)=\mathbb{P}(y_{x_{1},\mathbf{Z}_{x_{0}}})-\mathbb{P}(y_{x_{0}}),$ (2)
where $\mathbf{Z}$ represents the set of mediator variables, and $\mathbb{P}(y_{x_{1},\mathbf{Z}_{x_{0}}})$ is the probability of $Y=y$ if $X$ is set to $x_{1}$ while the mediators are set to values they would take under $X=x_{0}$.
Natural Indirect Effect (NIE) (Pearl, 2001): The NIE captures the influence of $X$ on $Y$ through mediators. It is given by:
$NIE_{x_{1},x_{0}}(y)=\mathbb{P}(y_{x_{0},\mathbf{Z}_{x_{1}}})-\mathbb{P}(y_{x_{0}}),$ (3)
where $\mathbb{P}(y_{x_{0},\mathbf{Z}_{x_{1}}})$ represents the probability of $Y=y$ when $X=x_{0}$ but mediators take values they would under $X=x_{1}$.
Path-Specific Effect (PSE) (Pearl, 2009b; Chiappa, 2019; Wu et al., 2019): The PSE isolates the causal effect of $X$ on $Y$ transmitted through a specific path or set of paths $\pi$. Formally, it is defined as:
$PSE^{\pi}_{x_{1},x_{0}}(y)=\mathbb{P}(y_{x_{1}\mid_{\pi},x_{0}\mid_{\overline{\pi}}})-\mathbb{P}(y_{x_{0}}),$ (4)
where $\mathbb{P}(y_{x_{1}\mid_{\pi},x_{0}\mid_{\overline{\pi}}})$ is the probability of $Y=y$ if $X=x_{1}$ along path $\pi$, while other paths ($\overline{\pi}$) remain unaffected by the intervention.
C.2.3 No Unresolved Discrimination
No unresolved discrimination (Kilbertus et al., 2017) requires that any causal effect of a sensitive attribute $X$ on an outcome $Y$ occurs only through resolving (explanatory) variables. A resolving variable, such as education level, reflects a non-discriminatory influence of $X$ on $Y$. The criterion prohibits direct and proxy effects of $X$ on $Y$.
C.2.4 No Proxy Discrimination
No proxy discrimination (Kilbertus et al., 2017) ensures that decisions are not influenced by variables $R$ that act as proxies for sensitive attributes $X$. Proxy discrimination is absent if:
$\mathbb{P}(Y\mid do(R=r))=\mathbb{P}(Y\mid do(R=r^{\prime})),\quad\forall r,r^{\prime}\in\text{dom}(R).$ (5)
This guarantees that changes in $R$ do not affect the outcome $Y$ if $R$ is a proxy for $X$.
C.2.5 Counterfactual Fairness
Counterfactual fairness (Kusner et al., 2017) requires that the outcome $Y$ for an individual remains the same in both factual and counterfactual scenarios. Formally, counterfactual fairness holds if:
$\mathbb{P}(y_{x_{1}}\mid\mathbf{V}=\mathbf{v},X=x_{0})=\mathbb{P}(y_{x_{0}}\mid\mathbf{V}=\mathbf{v},X=x_{0}),$ (6)
where $\mathbf{V}$ represents all other variables in the causal graph. This definition ensures fairness at the individual level by requiring that the sensitive attribute $X$ does not influence $Y$ in any hypothetical scenario.
Appendix D Causality and ML
The use of causality in AI falls mainly into one of two categories. The first approach is to employ artificial intelligence to enhance the qualitative discovery and/or quantification of causal connections from the data. The second one is to use causal tools to improve Machine Learning (ML) predictions. Next, we elaborate on both of these methods to combine causality and ML.
D.1 ML for causality
Causal Discovery Most of the techniques for obtaining causal quantities rely on knowing the causal structure of the data. It was previously assumed to be provided by experts. Recent advances in causal discovery offer algorithmic tools for recovering causal graphs from observational data. The basis for causal discovery is the probabilistic and graphical concepts of causality (Dawid, 2010). Two main groups of causal discovery algorithms can be distinguished based on their attempt to identify conditional or unconditional (including pairwise) independencies in the distribution from which the observational data is generated. The first category includes constraints and score-based algorithms such as PC (Le et al., 2016), FCI (Spirtes et al., 2000), and GES (Hauser & Bühlmann, 2012). They usually produce a partially oriented causal graph. The second category consists of algorithms based on causal asymmetries such as LiNGAM (Shimizu et al., 2006), and PNL (Zhang & Hyvarinen, 2012). The algorithms based on Kolmogorov’s (algorithmic) complexity assume that if knowing the shortest compression of one variable does not reveal the shorter compression of the other, two variables are considered independent (Janzing & Schölkopf, 2010; Schölkopf et al., 2022). The summary of the principles and performance for pairwise causal discovery is provided by Mooij et al. (Mooij et al., 2016). If the assumptions of the algorithms are satisfied, they are capable of identifying a unique causal graph or a causal direction between the two variables.
ML Tools for Causal Inference Supervised or semi-supervised machine learning methods can be used to estimate causal quantities from the data or for variable selection in situations with a high number of covariates (Kreif & DiazOrdaz, 2019; Aoki & Ester, 2022). ML algorithms such as, for example, logistic regression, bagging, random forest, and others, can be beneficial in estimating propensity scores used to estimate causal effects in the potential outcome framework (Lee et al., 2010; Tu, 2019).
LLMs for Causal Discovery The recent advancements in large language models (LLMs) have inspired their use in causal discovery (Kıcıman et al., 2023; Kasetty et al., 2024; Vashishtha et al., 2023; AI4Science & Quantum, 2023; Abdulaal et al., 2023; Khatibi et al., 2024). Most of the above methods involve the refinement of the statistically inferred causal graph by LLM. However, emerging research shows that, LLMs excel at synthesizing vast amounts of heterogeneous knowledge, making them well-suited for tasks that require the integration of diverse datasets, such as constructing full causal graphs based on scientific literature in diverse domains (Sheth et al., 2024; Afonja et al., 2024).
D.2 Causality for ML
One of the main arguments that motivated the use of causality for machine learning is that causal modeling can lead to more invariant or robust models (Schölkopf et al., 2022). The problem of overfitting and vulnerability to a domain shift is a known problem in ML. It is intuitive that learning the correlation between two phenomena, for example, rain and umbrellas, will not help to predict rain in situations where people prefer raincoats instead of umbrellas. A causal understanding of phenomena is more general to multiple circumstances. Following Pearl, ”…we may as well view our unsatiated quest for understanding how data is generated or how things work as a quest to acquire the ability to make predictions under a wider range of circumstances, including circumstances in which things are taken apart, reconfigured, or undergo spontaneous change” (Pearl, 2009a). One of the methods to combine the ML model with the causal approach is to incorporate causal knowledge (usually in the form of a complete or partial causal graph) in the learning process (Berrevoets et al., 2023, 2022). Causal representation learning is an attempt to combine latent variables derived from unstructured data and causal structure to arrive at a more invariant or fair model (Schölkopf et al., 2021; Mitrovic et al., 2020; Schölkopf et al., 2022; Wang et al., 2022). The causal structure can also be used for feature selection, assuming that it is known. Models based on direct causes to predict the outcome are considered more robust (Tople et al., 2020).

### v4

@@ 1 Introduction
In recent years, machine learning (ML) has made remarkable strides, driving breakthroughs in natural language processing (Achiam et al., 2023), computer vision (Brooks et al., 2024), and decision-making systems (Jia et al., 2024). These advancements have led to widespread adoption across diverse domains, including healthcare (Singhal et al., 2025), finance (Lee et al., 2024), education (Modi et al., 2024), and social media (Bashiri & Kowsari, 2024), where ML models now play a crucial role in diagnostics, algorithmic trading, personalized learning, and content recommendation.
@@
Causality. One of the most influential causal frameworks is Pearl’s structural causal models (SCMs), which provide a systematic approach to reasoning about causality and integrating it into machine learning (Pearl, 2009). This framework defines causality as the relationship between the variables where a change in one variable (the cause) directly leads to a change in another variable (the effect). It establishes a directional and often mechanistic link, distinguishing relationships arising from mere correlations.
@@
Position. Despite significant advancements in research on individual dimensions of trustworthy ML such as fairness, privacy, and explainability—there is a notable lack of efforts to integrate these dimensions into a cohesive and unified framework. Each ethical principle addresses distinct challenges, yet their interplay often involves intricate trade-offs, particularly concerning model performance metrics such as accuracy. For example, mitigating fairness-related biases may require adjustments that compromise predictive precision, while enhancing explainability can impose constraints on model complexity. We argue that systematically addressing these trade-offs is a critical step toward developing AI systems that are both ethically sound and operationally efficient. While causality has been applied to address individual challenges such as fairness or interpretability, its potential to address the intersection of these challenges has largely been overlooked (see LABEL:app:relatedwork for a detailed review). In this paper, we argue that integrating causality into ML and foundation models offers a way to balance multiple competing objectives of trustworthy AI.
@@ 2.1 Causality for Trade-offs in Trustworthy ML
Fairness vs. Accuracy. Most of the statistical fairness literature focuses on improving fairness metrics while preserving accuracy as much as possible (Feldman et al., 2015; Wei & Niethammer, 2022; Wang et al., 2021). However, fairness often comes at the cost of reduced accuracy, as mitigating bias may require either obscuring predictive features that also contribute to discrimination or constraining model predictions within fairness-imposed boundaries (Pinzón et al., 2022; Cooper et al., 2021; Zliobaite, 2015; Zhao & Gordon, 2022).
@@
A compelling example comes from the COMPAS dataset, where Black defendants were more likely to be classified as high-risk for recidivism. Traditional statistical debiasing approaches treat race as a direct cause of the risk score, but a causal analysis reveals that increased recidivism risk is confounded by heightened policing in predominantly Black neighborhoods. By explicitly modeling this causal structure, fairness-enhancing interventions can adjust for the effect of over-policing, ensuring that predictions reflect true recidivism risk rather than biased enforcement patterns. This results in a more accurate and fairer risk assessment (Chiappa, 2019; Zafar et al., 2017).
Conflicting Notions of Fairness. Fairness in ML is often constrained by conflicting definitions and measurement approaches. Friedler et al. (2021) highlight the fundamental tension between the “what you see is what you get” and “we are all equal” worldviews—where the former accepts disparities based on observed merit, while the latter seeks to correct historical inequalities. Causal graphs can crisply formulate different notions of fairness (Nabi & Shpitser, 2018; Chen et al., 2024), thus enabling feasible mitigation via path-specific causal effects (Avin et al., 2005).
@@ 2.2 Integrating Causality into ML
All of the above forms a causal ML cycle (Figure 1) in which ML is enhanced by causal knowledge, controlled by causal tools, and finally contributes to enriching scientific knowledge. We include a supplementary introduction to causality and causal ML in LABEL:app:causalpreliminaries and LABEL:app:causalML.
@@ 3 Causality for Trustworthy Foundation Models
Foundation models, including state-of-the-art multimodal systems like Large Language Models (LLMs) and vision-language models, have demonstrated exceptional capabilities across diverse tasks (Achiam et al., 2023; Rohan et al., 2023; Radford et al., 2023; Brooks et al., 2024). However, their reliability remains a concern due to issues like spurious correlations, hallucinations, and unequal representation. Trade-offs and causality in trustworthy foundation models remain underexplored as compared to traditional ML. In this section, we explore the potential for causality to improve fairness, explainability, privacy, and robustness in foundation models following slightly different taxonomy than in the previous section due to their unique challenges.
@@ 3.1 Dimensions of Trustworthy Foundation Models
Fairness vs. Accuracy. Causal frameworks have become integral to fairness interventions in LLMs by identifying and mitigating pathways that lead to unfair predictions (Madhavan et al., 2023; Cotta & Maddison, 2024). Counterfactual fairness ensures that sensitive attributes (e.g., gender, race) do not causally influence outcomes. For example, in job recommendation systems, counterfactual fairness guarantees identical recommendations for equally qualified candidates regardless of their gender (Madhavan et al., 2023). Methods like causal disentanglement isolate sensitive features from output-relevant causal factors, ensuring that spurious correlations, such as gender biases in job roles, do not propagate through the model (Zhou et al., 2023; Chen et al., 2024). SCMs further enable fairness-aware fine-tuning by disentangling causal effects. However, striving for diversity has been shown to introduce non-factual output in text-to-image models. In early 2024, Google’s AI tool, Gemini, faced criticism for generating historically inaccurate images, such as depicting America’s Founding Fathers as Black individuals and Nazis as racially diverse (Vincent, 2024). Here, causality could help distinguish historically impossible scenarios from desirable diversity, ensuring both fairness and factual integrity in AI-generated content. Mode collapse is another foundation model-specific fairness issue where models generate overly generic outputs, reducing diversity and disproportionately omitting minority group representations. Causal modeling can potentially help preserve minority information by explicitly capturing causal relationships, preventing spurious correlations from erasing underrepresented patterns.
Robustness vs. Accuracy. Causal frameworks address robustness by training models to rely on invariant causal relationships while penalizing reliance on dataset-specific spurious features (Wu et al., 2024). For instance, instead of associating ”doctor” with ”male,” causal invariance enforces reliance on task-relevant features like medical terminology (Zhou et al., 2023). Causal regularization further discourages attention to non-causal patterns during inference achieving better accuracy and robustness.
@@ 3.2 Integrating Causality in Foundation Models
Pre-training: Causal Representation Learning. By disentangling causal factors from non-causal ones, models can learn representations that separate meaningful causal features from irrelevant associations. Techniques such as causal embedding methods (Rajendran et al., 2024; Jiang et al., 2024), which can use training data annotated with causal labels, can guide models in identifying and prioritizing true causal relationships. This has been shown to reduce reliance on spurious correlations, such as gender-biased occupational associations (Zhou et al., 2023).
@@
Post-training: Fine-tuning. Fine-tuning on datasets specifically designed to highlight causal reasoning (e.g., datasets emphasizing cause-effect linguistic patterns) ensures that models learn causal-invariant patterns. Further, counterfactual data samples can also improve the fine-tuning. Synthetic counterfactual examples improve the model’s robustness to spurious correlations, similar to pre-training, but with better sample size efficiency. Frameworks like DISCO (Chen et al., 2022) generate diverse counterfactuals during fine-tuning to enhance OOD generalization for downstream tasks. Causally Fair Language Models (CFL) (Madhavan et al., 2023) use SCM-based regularization to detoxify outputs or enforce demographically neutral generation during post-training.  Wang & Culotta (2020) use causal reasoning to separate genuine from spurious correlations by computing controlled direct effects, ensuring robust performance.